# A Second Main Theorem for Entire Curves Intersecting Three Conics

### v1

@@ Abstract.
We establish a Second Main Theorem for entire holomorphic curves $f:\mathbb{C}\to\mathbb{P}^{2}$ intersecting a generic configuration of three conics $\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}$ in the complex projective plane $\mathbb{P}^{2}$. Using invariant logarithmic $2$-jet differentials with negative twists, we prove the estimate
@@
Nevanlinna theory, Second Main Theorem, entire curves, jet differentials, Siu’s slanted vector fields, Demailly-Semple tower
@@ 1.1. Motivation
In the complex projective plane $\mathbb{P}^{2}$, the logarithmic Kobayashi conjecture posits that if $\mathcal{C}\subset\mathbb{P}^{2}$ is a generic algebraic curve of degree $d\geqslant 5$, then the complement $\mathbb{P}^{2}\setminus\mathcal{C}$ should be hyperbolic [Kobayashi1970]. Here, hyperbolicity means that $\mathbb{P}^{2}\setminus\mathcal{C}$ does not contain the image of any nonconstant holomorphic map $f:\mathbb{C}\rightarrow\mathbb{P}^{2}$. This conjecture generalizes the classical Little Picard Theorem on the hyperbolicity of $\mathbb{P}^{1}\setminus\{\text{3 points}\}$, and is supported by explicit examples constructed by Zaidenberg [Zaidenberg1988]. For generic curves $\mathcal{C}$ of sufficiently large degree in $\mathbb{P}^{2}$, the conjecture was confirmed by Siu and Yeung [siu_yeung1996] by constructing explicit logarithmic $2$-jet differentials with negative twists. Later, building on the work [Demailly-Elgoul2000], El Goul [Elgoul2003] improved the degree bound to $15$ for very generic $\mathcal{C}$. Combining this with Siu’s strategy of using slanted vector fields [Siu2004] and a technique of Păun [mihaipaun2008], Rousseau [Rousseau2009] further reduced the degree bound to $14$.
These results for high-degree curves naturally extend to configurations with multiple components. In the case where the curve $\mathcal{C}=\cup_{i=1}^{q}\mathcal{C}_{i}$ decomposes into the union of $q\geqslant 2$ generic algebraic curves $\mathcal{C}_{i}$ of degree $d_{i}$, the hyperbolicity of the complement $\mathbb{P}^{2}\setminus\mathcal{C}$ has been established with smaller degree bounds $d=\sum_{i=1}^{q}d_{i}$ in various works. For instance, a result in [babets1984] settles the case where $q\geqslant 5$. The case $(q=4,\sum_{i=1}^{4}d_{i}\geqslant 5)$ was treated in [DSW1995-AJM]. The configuration of three conics was considered in [Grauert1989, DSW1995-AJM, Francois-Duval-2001] by different methods. Especially, in [DSW1995-duke], the authors employed the symmetric differentials $1$-forms to deal with the case of two conics and one cubic. Rousseau [Rousseau2009] addressed the case $q=2$ under the assumptions $(d_{1}=1,d_{2}\geqslant 11)$, $(d_{1}=2,d_{2}\geqslant 8)$, $(d_{1}=3,d_{2}\geqslant 5)$, or $(d_{1}=4,d_{2}\geqslant 4)$. Notably, for $q\geqslant 3$, Noguchi, Winkelmann, and Yamanoi [NWA2007] established the algebraic degeneracy of every entire curve in $\mathbb{P}^{2}\setminus\mathcal{C}$ under the mild condition $\sum_{i=1}^{q}d_{i}\geqslant 4$. Their result also extends to the corresponding higher-dimensional setting.
@@
In the nonlinear case, analogous results (without truncation) were obtained by Eremenko and Sodin [Eremenko-Sodin1992] via pluripotential theory, and by Ru [Minru2009] through a synthesis of Cartan’s method and Diophantine geometry. These result also extends to the corresponding higher-dimensional setting. For further background and recent advances in Nevanlinna theory, see [Noguchi-Winkelmann2014, Ru21].
@@
In the case of $n=2$, the defect relation (1.2) might be viewed as the weakest, albeit nontrivial, form of a Second Main Theorem in the vein of (1.1) with noneffective $c<\infty$. The primary objective of this paper is to establish an effective Second Main Theorem, with truncation level $1$, for three generic conics with an explicit constant $c=5$.
@@ 1.2. Two Key Theorems
The pursuit of a Second Main Theorem encounters a profound ontological obstacle: the very nature of entire curves exhibits an inherent and wild flexibility in large-scale geometry, as evidenced by the exotic examples in [HX2021, Xie2024, WX2025, CFX2025]. Our strategy constrains this flexibility by leveraging additional algebro-geometric structures to extract the underlying rigidity of entire curves, building on the following two key results.
@@
$\omega\in H^{0}\big(\mathbb{P}^{2},E_{k,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)\quad(k,m,t\geqslant 1),$
@@
For the definition of logarithmic jet differential bundles $E_{k,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$, see Section 2. Theorem 1.2 can be seen as a quantitative refinement of the Fundamental Vanishing Theorem for entire curves (cf. [Green-Griffiths1980, pp. 55-59], [Siu-Yeung1996, Lemma 1.2] and [Demailly1997, Corollary 7.9]). Although the original statement in [Huynh-Vu-Xie-2017] is formulated for hypersurfaces in $\mathbb{P}^{n}$, the proof extends directly to the case of simple normal crossing divisors in projective manifolds (cf. [Ru21, Theorem A10.7.2]).
@@ 1.3. Our First Result
The first result of this paper is an application of Theorem 1.2 in the case $k=1$.
@@
$c>\frac{1}{(d-3)\delta_{1}},\quad\text{where }\quad{\delta_{1}=\frac{1}{2}-\frac{\sqrt{9(d-1)^{2}-12(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})}}{6(d-3)}}.$
Here, the numerical assumption on $d_{1},d_{2},d_{3}$ is made to guarantee that the Euler characteristics are positive:
@@
Since the answer to Question 1.5 is currently unknown, 111 By applying the approach developed in this work, we can verify the vanishing for $1\leqslant m\leqslant 5$ for generic three conics. the key assumption (1.4) cannot be verified a priori. This situation vividly illustrates the principle that “the lower the degree, the harder the hyperbolicity” [Duval2021, p. 3]. To circumvent this obstacle, we employ invariant logarithmic $2$-jet differentials with negative twists.
@@
The constant $5$ in this theorem is smaller than the constant ${\approx}11.74$ appearing in Theorem 1.3 for the case of two conics and one cubic. This may seem counterintuitive at first, since the larger total degree $2+2+3>2+2+2$ would generally suggest a smaller constant. The explanation lies in the method of proof: here we employ not only $1$-jet techniques but also the more refined $2$-jet approach. A detailed analysis using the latter is expected to yield sharper constants not only for the two-conics-and-one-cubic case but also for other configurations covered by Theorem 1.3.
We now outline the main ideas underlying the proof of Theorem 1.6. As in the proof of Theorem 1.3, we can assume that $f$ is algebraically nondegenerate.
@@
  • If $f^{*}\omega_{1}\not\equiv 0$ or $f^{*}\omega_{2}\not\equiv 0$, then Theorem 1.2 directly implies the desired Second Main Theorem type estimate.
  • If $f^{*}\omega_{1}\equiv 0$ and $f^{*}\omega_{2}\equiv 0$, then the “independence” of $\omega_{1}$ and $\omega_{2}$ ensures that the lifting $f_{[1]}(\mathbb{C})\subset X_{1}$ is algebraically degenerate, where $X_{1}$ is the first level of the logarithmic Demailly–Semple tower (see Subsection 2.2) for $(\mathbb{P}^{2},\mathcal{C}=\cup_{i=1}^{3}\mathcal{C}_{i})$. Hence McQuillan’s Theorem 1.1 applies.
@@
In practice, however, this conceptual framework faces substantial challenges. Controlling the common base locus of $\omega$ and $\omega^{\prime}$—essential for ensuring their independence—requires detailed geometric information that is typically accessible only through explicit constructions of negatively twisted logarithmic jet differentials. While such constructions are feasible when the curve $\mathcal{C}$ is of high degree, as demonstrated by Siu and Yeung [siu_yeung1996], the low-degree case presents significantly greater difficulties. To date, not a single such differential has been explicitly described for low-degree curves, due to the intricate algebraic complexity of jet differentials (cf. [Merker2008]). Moreover, the subsequent elimination step introduces further difficulties, rendering a direct approach effectively intractable.
@@
1.6. Roadmap for Theorem 1.6: Towards an Effective Finite Constant $c$
@@
  • If $f^{*}\omega_{2}\not\equiv 0$, we can derive a Second Main Theorem-type estimate; see Theorem 6.4. The arguments are analogous to those for Theorem 1.2 but are refined in a crucial way, using the structure of Wronskians (see Proposition 6.3). This refinement represents a new contribution of this paper.
  • If $f^{*}\omega_{2}\equiv 0$, then the simultaneous vanishing of $f^{*}\omega_{1}$ and $f^{*}\omega_{2}$ implies that the lifting $f_{[1]}(\mathbb{C})$ is algebraically degenerate in the first level $X_{1}$ of the logarithmic Demailly–Semple tower for $(\mathbb{P}^{2},\mathcal{C})$, and hence McQuillan’s Theorem 1.1 applies.
@@ 1.7. New Vanishing Results and Proof Strategy
While such vanishing results have long been recognized as fundamental in complex hyperbolicity theory, a proof has remained out of reach for lack of a viable method. The principal innovation of this paper is a novel methodological framework that resolves this impasse.
Our proof of Key Vanishing Lemma I proceeds in three steps, detailed in Subsection 7.2:
  • Algebraic Reduction: We begin by transforming the problem of constructing invariant logarithmic jet differentials into that of solving a specific, tractable linear system, leveraging the GAGA principle to apply algebraic techniques in an analytic setting. The central challenge then becomes establishing bounds on the exponents in the denominators of the associated coefficients. This challenge is addressed by a judicious choice of basis and an analysis of the transition matrices between different frames.
@@
To demonstrate the efficacy of our approach, we apply the systematic framework developed in this article to establish new vanishing lemmas about $2$-jet differentials, which in turn yield improved degree bounds for the Kobayashi hyperbolicity conjecture in complex dimension two. Specifically, we refine prior results by lowering Păun’s [mihaipaun2008] bound from $18$ to $16$ for generic surfaces in $\mathbb{P}^{3}$, and Rousseau’s [Rousseau2009] bound from $14$ to $12$ for complements of generic curves in $\mathbb{P}^{2}$. It should be noted that the proofs of these vanishing results require subtle and novel adaptations of our framework, and their application toward the improved bounds relies on other ingredients. Complete details will be presented in a forthcoming paper. 333 Moreover, our method shows promising potential for further improving these degree bounds to $15$ and $11$, respectively. The main obstacle lies in the substantial computational resources required to verify the necessary vanishing lemmas, which exceeds the capacity of standard computers and calls for parallel computing. The implementation is theoretically feasible but requires porting and optimizing our Maple code in C—a task we deem accessible but have not yet pursued.
1.8. Proof Blueprint of Theorem 1.6 with an Effective Constant $c=5$
@@
The proof (see Subsection 7.3) introduces two new ingredients: the utilization of Fermat-type conics to facilitate computation; and a modulo $p=5$ technique that leverages Maple to effectively perform Gaussian elimination, thus proving that the associated system of linear equations with integer coefficients has only the trivial solution.
1.9. A Challenge
To avoid the use of any Key Vanishing Lemma and directly obtain, for three generic conics $\mathcal{C}$, a constant
[…]
in the Second Main Theorem, it suffices to verify the following conjecture.
Conjecture 1.10.
Let $\mathcal{C}\subset\mathbb{P}^{2}$ be a configuration of three conics with simple normal crossings. For any $\epsilon>0$, there exists a nonzero negatively twisted invariant logarithmic $2$-jet differential
$\omega\in H^{0}\!\left(\mathbb{P}^{2},\;E_{2,m}T^{*}_{\mathbb{P}^{2}}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\right)$
with $\overline{\theta}_{2}(X,\mathcal{C})-\epsilon<\tfrac{t}{m}<\overline{\theta}_{2}(X,\mathcal{C})$, such that its zero locus $\{\omega=0\}\subset X_{2}$ is irreducible and reduced modulo $\Gamma_{2}$.
We expect that such a multiplicity-one property to remain valid in a more general context, in particular for generic members $\mathcal{C}$ of a sufficiently varying family.
1.10. Structure of the Paper
Section 2 provides necessary background. Section 3 presents detailed Riemann–Roch computations that form the core of the proof of Theorem 1.3. Section 4 advances to more sophisticated Riemann–Roch computations, specifically designed to establish the existence of negatively twisted invariant logarithmic $2$-jet differentials for three generic conics. Section 5 details the application of Siu’s slanted vector fields in Case 1. Section 6 then introduces a refinement of the Demailly–El Goul argument, adapted to handle Case 2. The core contributions of the paper are contained in Section 7, which presents the proofs of Key Vanishing Lemmas I and II. Finally, Appendix A provides the Maple code used to verify these key vanishing results. Appendix B presents an observation by Pengchao Wang.

### v2

@@ Abstract.
We establish a Second Main Theorem for entire holomorphic curves $f:\mathbb{C}\to\mathbb{P}^{2}$ intersecting a generic configuration of three conics $\mathcal{C}=\mathcal{C}_{1}+\mathcal{C}_{2}+\mathcal{C}_{3}$ in the complex projective plane $\mathbb{P}^{2}$. Using invariant logarithmic $2$-jet differentials with negative twists, we prove the estimate
@@
The key innovation of our approach is establishing new vanishing lemmas of the form
$H^{0}\bigl(\mathbb{P}^{2},\,E_{2,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\bigr)=0$
for specific pairs $(m,t)$, achieved by combining algebro-geometric arguments with computer-assisted computations through a mod-$p$ reduction technique. This yields a systematic method for proving vanishing results for negatively twisted jet differentials—a key component in complex hyperbolic geometry.
@@
Nevanlinna theory, Second Main Theorem, entire curves, invariant logarithmic $2$-jet differentials, Siu’s slanted vector fields, Demailly-Semple tower
@@ 1.1. Motivation
In the complex projective plane $\mathbb{P}^{2}$, the logarithmic Kobayashi conjecture asserts that if $\mathcal{C}\subset\mathbb{P}^{2}$ is a generic algebraic curve of degree $d\geqslant 5$, then the complement $\mathbb{P}^{2}\setminus\mathcal{C}$ should be hyperbolic [Kobayashi1970]. Various notions of hyperbolicity exist; here we use the most straightforward one: $\mathbb{P}^{2}\setminus\mathcal{C}$ is said to be hyperbolic if there is no nonconstant holomorphic map $f:\mathbb{C}\to\mathbb{P}^{2}\setminus\mathcal{C}$. This conjecture generalizes the classical Little Picard Theorem on the hyperbolicity of $\mathbb{P}^{1}\setminus\{\text{3 points}\}$, and is supported by explicit examples constructed by Zaidenberg [Zaidenberg1988]. For generic curves $\mathcal{C}$ of sufficiently large degree in $\mathbb{P}^{2}$, the conjecture was confirmed by Siu and Yeung [siu_yeung1996] by constructing explicit logarithmic $2$-jet differentials with negative twists. Later, building on the work [Demailly-Elgoul2000], El Goul [Elgoul2003] improved the degree bound to $15$ for very generic $\mathcal{C}$. Combining this with Siu’s strategy of using slanted vector fields [Siu2004] and a technique of Păun [mihaipaun2008], Rousseau [Rousseau2009] further reduced the degree bound to $14$.
These results for high-degree curves naturally extend to configurations with multiple components. In the case where the curve $\mathcal{C}=\cup_{i=1}^{q}\mathcal{C}_{i}$ decomposes into the union of $q\geqslant 2$ generic algebraic curves $\mathcal{C}_{i}$ of degree $d_{i}$, the hyperbolicity of the complement $\mathbb{P}^{2}\setminus\mathcal{C}$ has been established with smaller degree bounds $d=\sum_{i=1}^{q}d_{i}$ in various works. For instance, a result in [babets1984] settles the case where $q\geqslant 5$. The case $(q=4,\sum_{i=1}^{4}d_{i}\geqslant 5)$ was treated in [DSW1995-AJM]. The configuration of three conics was considered in [Grauert1989, DSW1995-AJM, Francois-Duval-2001] by different methods. Dethloff, Schumacher, and Wong [DSW1995-duke] employed the symmetric differentials $1$-forms to deal with the case of two conics and one cubic. Rousseau [Rousseau2009] addressed the case $q=2$ under the assumptions $(d_{1}=1,d_{2}\geqslant 11)$, $(d_{1}=2,d_{2}\geqslant 8)$, $(d_{1}=3,d_{2}\geqslant 5)$, or $(d_{1}=4,d_{2}\geqslant 4)$. Notably, for $q\geqslant 3$, Noguchi, Winkelmann, and Yamanoi [NWA2007] established the algebraic degeneracy of every entire curve in $\mathbb{P}^{2}\setminus\mathcal{C}$ under the mild condition $\sum_{i=1}^{q}d_{i}\geqslant 4$. Their result also extends to the corresponding higher-dimensional setting.
@@
In the nonlinear case, analogous results (without truncation) were obtained by Eremenko and Sodin [Eremenko-Sodin1992] via pluripotential theory, and by Ru [Minru2009] through a synthesis of Cartan’s method and Diophantine geometry. These results also extend to the corresponding higher-dimensional setting. For further background and recent advances in Nevanlinna theory, see [Noguchi-Winkelmann2014, Ru21].
@@
In the case of $n=2$, the defect relation (1.2) might be viewed as the weakest, albeit nontrivial, form of a Second Main Theorem in the vein of (1.1) with noneffective $c<\infty$. The primary objective of this article is to establish an effective Second Main Theorem, with truncation level $1$, for three generic conics with an explicit constant $c=5$.
@@ 1.2. Two Key Theorems
The pursuit of a Second Main Theorem encounters a fundamental obstacle: the very nature of entire curves exhibits an inherent and wild flexibility in large-scale geometry, as evidenced by the exotic examples in [HX2021, Xie2024, WX2025, CFX2025]. Our strategy constrains this flexibility by leveraging additional algebro-geometric structures to extract the underlying rigidity of entire curves, building on the following two key results.
@@
$\omega\in H^{0}\big(\mathbb{P}^{2},E_{k,m}^{GG}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\otimes\mathcal{O}_{\mathbb{P}^{2}}(-t)\big)\quad(k,m,t\geqslant 1),$
@@
Here, $E_{k,m}^{GG}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$ denotes the logarithmic Green–Griffiths $k$-jet bundle of weighted order $m$ (see Section 2). The use of such higher-order jet differentials in the study of complex hyperbolicity goes back to André Bloch [Bloch1926]. In the present work, we restrict attention to the “invariant” subbundle $E_{k,m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})\subset E_{k,m}^{GG}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathcal{C})$, originally introduced by Demailly [Demailly1997]. This subbundle possesses stronger positivity properties, making it particularly suitable for constructing sections; further details can be found in Section 2.
Theorem 1.2 can be seen as a quantitative refinement of the Fundamental Vanishing Theorem for entire curves (cf. [Green-Griffiths1980, pp. 55-59], [Siu-Yeung1996, Lemma 1.2] and [Demailly1997, Corollary 7.9]). Although the original statement in [Huynh-Vu-Xie-2017] is formulated for hypersurfaces in $\mathbb{P}^{n}$, the proof extends directly to the case of simple normal crossing divisors in projective manifolds (cf. [Ru21, Theorem A10.7.2]).
@@ 1.3. Our First Result
The first result of this article is an application of Theorem 1.2 in the case $k=1$.
@@
$c>\frac{1}{(d-3)\delta_{1}},\quad\text{where }\quad{\delta_{1}=\tfrac{1}{2}-\tfrac{\sqrt{9(d-1)^{2}-12(d_{1}d_{2}+d_{2}d_{3}+d_{3}d_{1})}}{6(d-3)}}.$
Here, the numerical assumption on $d_{1},d_{2},d_{3}$ is made to guarantee that the Euler characteristics of the logarithmic symmetric differential bundles $S^{m}T_{\mathbb{P}^{2}}^{*}(\mathrm{log}\,\mathscr{C})$ are positive:
@@
Since the answer to Question 1.5 remains unknown, [^1] Using the approach developed in this work, we can verify the vanishing for $1\leqslant m\leqslant 7$ for generic three conics. the key assumption (1.4) cannot be verified a priori. This situation illustrates the principle that “the lower the degree, the harder the hyperbolicity” [Duval2021, p. 3]. To circumvent this obstacle, we employ invariant logarithmic $2$-jet differentials with negative twists. The main result of this article is as follows.
@@
The constant $c=5$ in this theorem is smaller than the constant ${\approx}11.74$ appearing in Theorem 1.3 for the case of two conics and one cubic. This may seem counterintuitive at first, since the smaller total degree $2+2+2<2+2+3$ would generally suggest a larger constant. The explanation lies in the method of proof: here we employ not only $1$-jet technique but also the more refined $2$-jet approach. A detailed analysis using the latter is expected to yield sharper constants not only for the two-conics-and-one-cubic case but also for other configurations covered by Theorem 1.3.
Remark 1.8.
An unexpected outcome of our method is that, as a byproduct, it allows us to explicitly identify certain configurations of three Fermat-type conics $\mathcal{C}$ satisfying the Second Main Theorem type estimate in (1.5) via an alternative “direct” argument. For instance, consider the three conics $\mathcal{C}$ defined by the quadratic polynomials:
$2Z_{0}^{2}+Z_{1}^{2}+Z_{2}^{2},\qquad Z_{0}^{2}+2Z_{1}^{2}+Z_{2}^{2},\qquad Z_{0}^{2}+Z_{1}^{2}+2Z_{2}^{2}.$
Further details can be found in Remark 7.5. This also explains why we are satisfied with the constant $5$ in the Main Theorem, rather than attempting to lower it slightly by proving additional vanishing results—which, while manageable, are not very interesting.
We now outline the main ideas underlying the proof of Theorem 1.6 for a finite constant $c<\infty$, before refining the argument to achieve the sharper goal $c=5$. As in the proof of Theorem 1.3, we can assume that $f$ is algebraically nondegenerate.
@@
  • If $f^{*}\omega_{1}\not\equiv 0$ or $f^{*}\omega_{2}\not\equiv 0$, then Theorem 1.2 directly implies a desired Second Main Theorem type estimate.
  • If $f^{*}\omega_{1}\equiv 0$ and $f^{*}\omega_{2}\equiv 0$, then the “independence” of $\omega_{1}$ and $\omega_{2}$ ensures that the lifting $f_{[1]}(\mathbb{C})\subset X_{1}$ is algebraically degenerate, where $X_{1}$ is the first level of the logarithmic Demailly-Semple tower (see Subsection 2.2) for $(\mathbb{P}^{2},\mathcal{C})$. Hence McQuillan’s Theorem 1.1 applies.
@@
In practice, however, this conceptual framework faces substantial challenges. Controlling the common base locus of $\omega_{1}$ and $\omega_{2}$—essential for ensuring their independence—requires detailed geometric information that is typically accessible only through explicit constructions of negatively twisted logarithmic jet differentials. While such constructions are feasible when the curve $\mathcal{C}$ is of high degree, as demonstrated by Siu and Yeung [siu_yeung1996], the low-degree case presents significantly greater difficulties. To date, not a single such differential has been explicitly described for low-degree curves, due to the intricate algebraic complexity of jet differentials (cf. [Merker2008]). Moreover, the subsequent elimination step introduces further difficulties, rendering a direct approach effectively intractable.
@@
1.6. Roadmap for Theorem 1.6: Towards an Effective Finite Constant $c<\infty$
@@
  • If $f^{*}\omega_{2}\not\equiv 0$, we can derive a Second Main Theorem-type estimate; see Theorem 6.4. The arguments are analogous to those for Theorem 1.2 but are refined in a crucial way, using the structure of Wronskians (see Proposition 6.3). This refinement represents a new contribution of this article.
  • If $f^{*}\omega_{2}\equiv 0$, then the simultaneous vanishing of $f^{*}\omega_{1}$ and $f^{*}\omega_{2}$ implies that the lifting $f_{[1]}(\mathbb{C})$ is algebraically degenerate in the first level $X_{1}$ of the logarithmic Demailly-Semple tower for $(\mathbb{P}^{2},\mathcal{C})$, and hence McQuillan’s Theorem 1.1 applies.
@@ 1.7. New Vanishing Results and Proof Strategy
While such vanishing results have long been recognized as fundamental in complex hyperbolicity theory, a proof has remained out of reach for lack of a viable method. The principal innovation of this article is a novel methodological framework that resolves this impasse.
Our proof of Key Vanishing Lemma I proceeds in three steps, presented in Subsection 7.2:
  • Algebraic Reduction: We begin by transforming the problem of constructing invariant logarithmic jet differentials into that of solving a specific, tractable linear system, leveraging the gaga principle to apply algebraic techniques in an analytic setting. The central challenge then becomes establishing bounds on the exponents in the denominators of the associated coefficients. This challenge is addressed by a careful choice of basis and an analysis of the transition matrices between different frames.
@@
To demonstrate the efficacy of our approach, we apply the systematic framework developed in this article to establish new vanishing results about $2$-jet differentials, which in turn yield improved degree bounds for the Kobayashi hyperbolicity conjecture in complex dimension two. Specifically, we refine prior results by lowering Păun’s [mihaipaun2008] bound from $18$ to $16$ for very generic surfaces in $\mathbb{P}^{3}$, and Rousseau’s [Rousseau2009] bound from $14$ to $12$ for complements of generic curves in $\mathbb{P}^{2}$. Complete details will be presented in a forthcoming paper.
1.8. Proof Blueprint of Theorem 1.6 with $c=5$
@@
The proof (see Subsection 7.3) introduces two new ingredients: the utilization of Fermat-type conics to facilitate computation; and a modulo $p=5$ technique that leverages computer effectively perform Gaussian elimination, thus proving that the associated system of linear equations with integer coefficients has only the trivial solution.
1.9. Structure of the article
Section 2 provides necessary background. Section 3 presents detailed Riemann–Roch computations that form the core of the proof of Theorem 1.3. Section 4 advances to more sophisticated Riemann–Roch computations, specifically designed to establish the existence of negatively twisted invariant logarithmic $2$-jet differentials for three generic conics. Section 5 details the application of Siu’s slanted vector fields in Case 1. Section 6 then introduces a refinement of the Demailly–El Goul argument, adapted to handle Case 2. The main innovation of the article is contained in Section 7, where we overcome fundamental obstructions in our low-degree setting by establishing Key Vanishing Lemmas. Finally, Appendix A provides the Maple code used to verify these vanishing results, and Appendix B presents an observation by Pengchao Wang.